FIFO Muskoge Company

FIFO Method, Two-Department Analysis Muskoge Company uses a process-costing system. The company manufactures a product that iCosts: Transferred in ? $56,250 $39,550 Direct materials Conversion costs Percentage of completion: $106,500 $136,500 BeginniRequired: 1. Using the FIFO method, prepare the following for the Molding Department: a. A physical flow schedule Muskoge Comb. An equivalent units calculation Total Equivalent Units Direct Materials 25,000 Conversion Costs 30,000 c. Calculation of ud. Cost of ending work in process and cost of goods transferred out. Cost of ending work in process: $ 25,050 x Cost of goods3. Repeat Requirements 1 and 2 for the Assembly Department. a. A physical flow schedule Muskoge Company Assembly Department Pb. An equivalent units calculation Total Equivalent Units Direct Materials 35,000 Conversion Costs 32,500 ✓ Transferred In 30
Answer #1

I. For Molding Department

Beginning inventory - 10,000 units which is 40% complete [i.e. 100% material (since material is applied in the beginning) and 40% conversion cost] at a cost of $ 35,800 ($ 22,000 for material and $ 13,800 for conversion).

Computing ending inventory -

Ending inventory = Beginning inventory + Units started - Units transferred

= 10,000 + 25,000 - 30,000

= 5,000 units which is 80% complete as given in the question

Reconciliation of units

Beginning WIP + Units started = Units completed from beginning WIP + Units started and completed + Ending inventory

10,000 (given) + 25,000 (given) = 10,000 (since FIFO method) + 20,000 (total completed units is 30,000 out of which 10,000 are from beginning inventory) + 5,000 (as computed above)

35,000 units = 35,000 units

Material and conversion process incurred during the relevant period in equivalent units

Material

Conversion cost

Beginning inventory (10,000)

Nil (0%)

6,000 (60%)

Started and completed (20,000)

20,000 (100%)

20,000 (100%)

Ending inventory (5,000)

5,000 (100%)

4,000 (80%)

Total

25,000

30,000

Cost incurred during the period (in $)

56,250

106,500

Cost per unit

$ 2.25 (56,250/25,000)

$ 3.55 (106,500/ 30,000)

Hence, total cost per unit is $ 5.80 per unit ($2.25 + $3.55).

Computation of cost

Beginning inventory

Started and completed

Ending inventory

Opening cost (A)

35,800 (given)

-

-

Material cost (B)

Nil

45,000 (20,000 * 2.25)

11,250 (5,000 * 2.25)

Conversion cost (C)

21,300 (6,000 * 3.55)

71,000 (20,000 * 3.55)

14,200 (4,000 * 3.55)

Total cost incurred during the period (D = B + C)

21,300

116,000

25,450

Total cost (E = A + D)

57,100

116,000

25,450

Cost incurred during the period = 162,750 (21,300 + 116,000 + 25,450)

Cost of units transferred = 173,100 (57,100 + 116,000)

Cost of ending inventory = 25,450

Total cost accounted for = 198,550 (173,100 + 25,450)

II. For assembly Department

Beginning inventory - 8,000 units which is 50% complete [i.e. Nil material (since material is applied at the end) and 50% conversion cost] at a cost of $61,100 ($45,400 for material and $ 16,700 for conversion).

Computing ending inventory -

Ending inventory = Beginning inventory + Units introduced - Units completed

= 8,000 + 30,000 - 35,000

= 3,000 units which is 50% complete as given in the question

Reconciliation of units

Beginning WIP + Units introduced = Units completed from beginning WIP + Units started and completed + Ending inventory

8,000 (given) + 30,000 (given) = 8,000 (since FIFO method) + 27,000 (total completed units is 35,000 out of which 8,000 are from beginning inventory) + 3,000 (as computed above)

38,000 units = 38,000 units

Material and conversion process incurred during the relevant period in equivalent units

Material

Conversion cost

Beginning inventory (8,000)

8,000 (100%)

4,000 (50%)

Started and completed (27,000)

27,000 (100%)

27,000 (100%)

Ending inventory (3,000)

Nil (0%)

1,500 (50%)

Total

35,000

32,500

Cost incurred during the period (in $)

39,550

136,500

Cost per unit

$ 1.13 (39,550/35,000)

$ 4.20 (136,500/ 32,500)

Hence, total cost per unit is $ 5.33 per unit ($1.13 + $4.20).

Computation of cost

Beginning inventory

Started and completed

Ending inventory

Opening cost (A)

62,100 (given)

-

-

Material cost (B)

9,040 (8,000 * 1.13)

30,510 (27,000 * 1.13)

Nil (Nil * 1.13)

Conversion cost (C)

16,800 (4,000 * 4.20)

113,400 (27,000 * 4.20)

6,300 (1,500 * 4.20)

Total cost incurred during the period (D = B + C)

25,840

143,910

6,300

Total cost (E = A + D)

87,940

143,910

6,300

Cost incurred during the period = 176,050 (25,840 + 143,910 + 6,300)

Cost of units transferred out = 231,850 (87,940 + 143,910)

Cost of ending inventory = 6,300

Total cost accounted for = 238,150 (231,859 + 6,300)

Cost reconciliation

Cost of beginning inventory                        =             62,100

Cost incurred during the period                 =             1,76,050

Total                                                                      =             238,150

Cost of units transferred out                      =             231,850

Ending WIP                                                         =             6,300

Total                                                                      =             238,150

Lastly, the cost of 30,000 units transferred in Assembly division from Molding division is $ 173,100 (as computed in Part I)